OpenStudy (anonymous):

The manager of a large apartment complex knows from experience that 120 units will be occupied if the rent is 400 dollars per month. A market survey suggests that, on the average, one additional unit will remain vacant for each 4 dollar increase in rent. Similarly, one additional unit will be occupied for each 4 dollar decrease in rent. What rent should the manager charge to maximize revenue?

OpenStudy (anonymous):

I just need to know how to set it up. Like the formula to use.

ganeshie8 (ganeshie8):

Revenue = price*sales

OpenStudy (anonymous):

does that work for any problem like this?

ganeshie8 (ganeshie8):

say, p = prince of each unit s = total number of units sold/occupied

ganeshie8 (ganeshie8):

yes, revenue equation remains same. it will not change :)

ganeshie8 (ganeshie8):

R = p*s

OpenStudy (anonymous):

For example, "A manufacture has been selling 1450 television sets a week at $390 each. A market survey indicates that for each $15 rebate offered to a buyer, the number of sets sold will increase by 150 per week. a) Find the function representing the demand p(x), where x is the number of the television sets sold per week and p(x) is the corresponding price."

ganeshie8 (ganeshie8):

lets finish our original problem first ?

OpenStudy (anonymous):

Ok but I think I got the majority of it

ganeshie8 (ganeshie8):

oh ok good :) wat did u get for rent that maximizes the revenue ?

OpenStudy (anonymous):

q(p) = 120 - (p-400) = 520 - p ?

ganeshie8 (ganeshie8):

whats q(p) ?

OpenStudy (anonymous):

Sorry it's just a letter I like to use lol... Is the answer 440?

ganeshie8 (ganeshie8):

thats right :- we're given : ds/dp = -1/4 and we know : R = p*s differentiate with respect to p both sides dR/dp = p*ds/dp + s = -p/4 + s set it equal to 0, u wud get p = 4s solving u wud get, p = 440, s = 110

OpenStudy (anonymous):

so then I would do kinda the same thing for the other one I showed you?

ganeshie8 (ganeshie8):

the other one asks for DEMAND. DEMAND is not same as REVENUE ok

OpenStudy (anonymous):

so what formula would I use?

ganeshie8 (ganeshie8):

For example, "A manufacture has been selling 1450 television sets a week at $390 each. A market survey indicates that for each $15 rebate offered to a buyer, the number of sets sold will increase by 150 per week. a) Find the function representing the demand p(x), where x is the number of the television sets sold per week and p(x) is the corresponding price."

ganeshie8 (ganeshie8):

for part a), you just need to create the price as a function of sales

ganeshie8 (ganeshie8):

say price of tvset = p number of television sets sold = x

ganeshie8 (ganeshie8):

p = f(x)

ganeshie8 (ganeshie8):

\(p(x) = 390 + \frac{15}{150}(1450-x) \)

ganeshie8 (ganeshie8):

thats the equation of price, see if that makes sense

OpenStudy (anonymous):

Yep thats what I have written down :) P=-.1x+535

ganeshie8 (ganeshie8):

Looks good :) thats all we're done for part a

OpenStudy (anonymous):

Ok thank you. Final tomorrow! :O

ganeshie8 (ganeshie8):

good luck !

ganeshie8 (ganeshie8):

once u get the equation, its always a good idea to make sure u got it right for ex :- u said u got P=-.1x+535

ganeshie8 (ganeshie8):

to check if its correct, simply put x = 1450 u should get the corresponding price 390

ganeshie8 (ganeshie8):

P=-.1x+535 put x = 1450 P = -.1*1450 + 535 = -145 + 535 = 390 so our equation for p(x) is \(\large \color{red}{\checkmark}\)

ganeshie8 (ganeshie8):

just a paranoid check ok :)

OpenStudy (anonymous):

It's a multiple choice final so that will help Thanks! :)